OpenStudy (anonymous):

3(z + 1) + 4(z – 3) = 5

OpenStudy (anonymous):

3z + 3 + 4z -12 = 5 add like terms: 7z - 9 = 5

OpenStudy (anonymous):

3z+3+4z-12=5 7z-9=5 7z=5+9 7z=14 z=2

OpenStudy (hihi67):

3*z=3z 3*1=3 + 4*z=4z 4*-3=-12 sooo 3z+3+4z-12=5

OpenStudy (hihi67):

and then..

OpenStudy (hihi67):

3z+4z =7z 3-12=-9 sooo 7z-9=5

OpenStudy (hihi67):

then.. 7z-9=5 +9 7z=14 /7 z=2

OpenStudy (anonymous):

thank you